How many pounds of 20% copper alloy must be mixed with 600 lb. of a 30% copper alloy to make a 27.5% copper alloy? (must show alg)

Answer by checkley77(12844)

.2x+600*.3=.275(600+x)
.2x+180=165+.275x
.2x-.275x=165-180
-.075=-15
x=-15/-.075
x=200 lbs of 20% copper is needed.
proof:
.2*200+.3*600=.275*800
40+180=220
220=220